OpenStudy (anonymous):

Find the area of the region between y=x^2+6x and the x axis for the following intervals. a.) -6 to 0 b.) -7 to 0 c.) -6 to 2

OpenStudy (anonymous):

so the x axis happens to be when y=0

OpenStudy (anonymous):

so the region is between y=x^2+6x and y=0

OpenStudy (anonymous):

first you have to draw the graph

OpenStudy (anonymous):

or at least figure out which is always on top

OpenStudy (anonymous):

if plug in your intervals you'll find that all the intervals start at intersections or where y=x^2+6x is above y=0

OpenStudy (anonymous):

\[\int f(x)-g(x)dx\] where fx is alway the above function and g(x) is the lower function

OpenStudy (anonymous):

so your integral will always be \[\int_{c}{}x^2+6xdx\] where c is your intervals

OpenStudy (anonymous):

can you solve this?

OpenStudy (anonymous):

no, still new to this

OpenStudy (anonymous):

Alright well can you watch this video and let me know if this helps http://patrickjmt.com/finding-areas-between-curves/

OpenStudy (anonymous):

this is basically what i explained earlier. Do you know how to find anti derivatives of polynomials?

OpenStudy (anonymous):

yes i am learning to do anti derivatives

OpenStudy (anonymous):

alright well watch the above video when you're done let me know we'll go on

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

outkast what the highest level math you took?

OpenStudy (anonymous):

I'm currently taking Differential Equations